OpenStudy (gabylovesyou):

Solve the equation 4/5y + 2 = 6 for y. Please show your work and solution in the space provided.

OpenStudy (anonymous):

4/5y + 2 = 6 subtract 2 from both sides 4/5y = 4 then multiply both sides by 5/4 to get rid of 4/5 y=4 X 5/4 which equals... 20/4 which equals y=5